A solution is made by dissolving 125 g of sodium chloride in 75.0 kg of water what is the percent by mass?


1
Expert's answer
2022-03-02T03:10:06-0500

Total mass "=75+125=200g"



"\frac{125}{200}×100\%=62.5\%"